OpenStudy (anonymous):

MEDAL, FAN, AND TESTIMONIAL In the manufacturing process of sulfuric acid, sulfur dioxide is reacted with oxygen to produce sulfur trioxide. Using the equation, 2SO2 (g) + O2 -> 2SO3 (g), if 128 g of sulfur dioxide is given the opportunity to react with an excess of oxygen, but only produces 144 g of sulfur trioxide, what is the percent yield of this reaction? 29.24% 58.48% 90.0% 100.0%

OpenStudy (anonymous):

Molar mass of SO2 = 64g So 2 moles of SO2 is being used. 2 moles of SO2 used => 2 moles of SO3 produced. Mass of SO3 produced = 144g Molar mass of SO3 = 80g => No of moles of SO3 = 144/80 = 1.8 % yield = 1.8*100/2 = 90%

OpenStudy (anonymous):

90%